OpenStudy (anonymous):

Solve the equation for y: 4x + 5(y + 3) = 10.

OpenStudy (anonymous):

4x + 5y + 15 = 10 Then. 4x + 5y = -5 5y = -5 - 4x y = -5/5 - 4x/5 y = -1 - 4x/5